OpenStudy (anonymous):

The gas gauge on a van initially read 1/8 full. When 6 gallons were added to the tank, the gauge read 1/2 full. How many more gallons are needed to fill the tank?

OpenStudy (aaronq):

so 6 gallons are equal to \(\dfrac{1}{2}-\dfrac{1}{8}\) or otherwise \(\dfrac{4}{8}-\dfrac{1}{8}=\dfrac{3}{8}\) so 2 gallons is equal to \(\dfrac{1}{8}\) of the tank. we need \(\dfrac{1}{2}\) or \(\dfrac{4}{8}\) the tank

OpenStudy (anonymous):

thank you for laying it out!!

OpenStudy (aaronq):

no problem!